PLEASE ANSWER ALL OF THE FOLLOWING. WORKDOES NOT NEED TO BE SHOWN OR EXPLAINED, JUST...


1.The following table shows the worldwide smartphone shipments of 2016.

Vendor2016 Shipments (millions of units)
Samsung309.4
Apple215.4
Huawei138.8
OPPO84.6
Vivo71.9
Others668.1

The market share of Apple's iphone in 2016 is ( %) (answer format: XX.X, round to one digit after decimal).

2. To show the market share of the top five smartphone manufacturers, we normally use ______

pie chart

histogram

side by side bar chart

ogive graph

Pareto chart
Answer #1

1) Market share of apple = (215.4/total)*100% = (215.4/1488.2)*100% = 14.4739 %

2) Pie chart